### v2

# Quantum Incompatibility in Parallel versus Antiparallel Spins

Abstract
We explore the joint measurability of incompatible qubit observables on ensembles of parallel and antiparallel spin-½ pairs. In parallel configuration, both spins are prepared in the same state, whereas in antiparallel case, each spin is paired with its flipped counterpart. We show that the antiparallel configuration uniquely enables the exact simultaneous prediction of three mutually orthogonal spin components—an advantage not achievable with parallel states. Extending beyond three observables, we examine joint measurability for larger sets of spin measurements and further generalize our analysis to state configurations beyond the parallel and antiparallel cases. As we show, our results reveal a deep connection to the ‘mean King retrodiction task’ proposed by Vaidman, Aharonov, and Albert, and have implications for a cryptographic protocol introduced by Jeffrey Bub. We further demonstrate how the enhanced compatibility in the antiparallel configuration can facilitate efficient estimation of unknown measurement devices. Finally, we discuss prospects for experimentally realizing the enhanced measurement compatibility in antiparallel configuration by analyzing the effect on finite sub-ensembles of states.
Introduction.– Bohr’s complementarity principle, a cornerstone of quantum theory, imposes fundamental limitation on simultaneous measurement of certain observables [1]. This is famously exemplified by the trade-off between path information and interference visibility in the double-slit experiment [2, 3, 4], as well as by the impossibility of jointly measuring non-commuting observables such as position and momentum, or spin components along different axes [5, 6, 7]. Development of generalized measurements, formalized via positive operator-valued measures (POVMs) [8], refined this understanding by demonstrating that incompatible observables can, in fact, be jointly measured—albeit with an inherent degree of fuzziness or imprecision [9, 10, 11, 12]. Lately, measurement incompatibility has been shown to be intimately connected to other nonclassical phenomena, such as Bell nonlocality and Einstein-Podolsky-Rosen steering [13, 14, 15, 16, 17, 18, 19, 20, 21, 22]. Measurement incompatibility also plays a pivotal role in several quantum protocols [12], and thus motivates extensive investigations into the structure of incompatibility, including settings involving multiple copies of quantum systems [23].
In this Letter, we investigate whether the state configuration in a multi-copy setup offers any advantage in joint measurability. Focusing on the two-copy case, we compare two configurations: the parallel configuration of two identically prepared spin-½ particles (qubits) and the antiparallel configuration of a qubit paired with its spin-flipped counterpart. We demonstrate that a measurement device operating on antiparallel spin pairs can reproduce the statistics of spin measurements along the $x$, $y$, and $z$ directions—outperforming what is possible in the parallel case. Furthermore, we show that perfect joint measurability extends to more than three incompatible spin observables in the antiparallel setting. Moving beyond these two canonical cases, we consider more general configuration of the form $\rho_{\vec{m}}\otimes\Lambda(\rho_{\vec{m}})$, where $\Lambda$ is a positive trace-preserving (PTP) map [24, 25, 26, 27]. Our construction also reveals a deep connection to other foundational concepts, such as the mean King retrodiction task [28], and enables improved robustness in the quantum key distribution protocol proposed by Bub [29]. As we explore the perfect joint measurability of mutually unbiased qubit bases in the antiparallel configuration suggests a resource-efficient strategy for estimating unknown measurement devices. Finally, we discuss potential pathways toward experimental realization of our scheme by analyzing the notion of measurement compatibility on finite sub-ensembles of states.
Measurement compatibility of unsharp spin observables.– The state of a spin-½ particle is described by a qubit density operator $\rho_{\vec{m}}=\tfrac{1}{2}(\mathbf{1}+\vec{m}\cdot\vec{\sigma})\in\mathcal{D}(\mathbb{C}^{2})$, where $\vec{m}\in\mathbb{R}^{3}$ with $|\vec{m}|\leq 1$, $\vec{m}\cdot\vec{\sigma}:=m_{x}\sigma_{x}+m_{y}\sigma_{y}+m_{z}\sigma_{z}$, the $\sigma$’s are the Pauli operators, and $\mathbf{1}$ is the identity operator. The state is pure when $|\vec{m}|=1$, and mixed otherwise. An unsharp spin measurement along $\hat{n}$ is defined as $\sigma_{\hat{n}}(\lambda)\equiv\{\mathrm{P}^{a}_{\hat{n}}(\lambda)=\tfrac{1}{2}(\mathbf{1}+\lambda\,a\,\hat{n}\cdot\vec{\sigma})\penalty 10000\ |\penalty 10000\ a=\pm 1\}$, where $\lambda\in[0,1]$ quantifies the sharpness of the measurement: $\lambda=1$ corresponds to a projective (sharp) measurement, while $\lambda=0$ corresponds to a completely uninformative (i.e., random-guess) measurement [30]. More generally, a positive operator-valued measure (POVM) $\mathcal{M}$ is a collection of effects summing to the identity, i.e., $\mathcal{M}\equiv\{\pi_{k}\in\mathcal{E}(\mathbb{C}^{2})\penalty 10000\ |\penalty 10000\ \sum_{k}\pi_{k}=\mathbf{1}\}$ [31]. This framework allows us to define joint measurability of spin observables.
Definition 1.
[Busch et al. [10]] A set of spin observables $S\equiv\{\sigma_{\hat{n}_{r}}(\lambda)\}_{r}$ is said to be jointly measurable if there exists a POVM $\mathcal{G}\equiv\{\pi_{{\bf a}}\}_{{\bf a}}$ such that $\mathrm{P}^{a_{r}}_{\hat{n}_{r}}(\lambda)=\sum_{{\bf a}\setminus a_{r}}\pi_{{\bf a}},\penalty 10000\ \forall\penalty 10000\ r,a_{r}$.
Here, ${\bf a}:=[a_{1},\dots,a_{\mathrm{N}}]\in\{+1,-1\}^{\mathrm{N}}$, where $\mathrm{N}$ is the cardinality of the set $S$, and ${\bf a}\setminus a_{r}$ denotes the summation over all components except $a_{r}$. While sharp spin observables along distinct directions are not jointly measurable, their unsharp versions can be, for suitable values of $\lambda$. For instance, spin observables along the $x$ and $y$ directions are jointly measurable when $\lambda\leq 1/\sqrt{2}$, and observables along $x$, $y$, and $z$ directions are jointly measurable when $\lambda\leq 1/\sqrt{3}$ [30]. For more recent developments, see [32, 33, 34].
Measurement compatibility in multi-copy setting.– More recently, Carmeli et al. [23] have investigated measurement compatibility in a multi-copy setting, where the experimenter has access to multiple copies of a quantum state in each measurement run.
Definition 2.
[Carmeli et al. [23]] A set $S$ of $N$ spin observables is $\mathrm{K}$-copy jointly measurable if there exists a POVM $\mathcal{G}\equiv\{\pi_{{\bf a}}\in\mathcal{P}((\mathbb{C}^{2})^{\otimes\mathrm{K}})\penalty 10000\ |\penalty 10000\ \sum_{{\bf a}}\pi_{{\bf a}}=\mathbf{1}^{\otimes\mathrm{K}}\}$ such that, for all states $\rho_{\vec{m}}$ and all $r,\penalty 10000\ a_{r}$, $\mathrm{Tr}[\rho_{\vec{m}}\,\mathrm{P}^{a_{r}}_{\hat{n}_{r}}(\lambda)]=\sum_{{\bf a}\setminus a_{r}}\mathrm{Tr}[\rho_{\vec{m}}^{\otimes\mathrm{K}}\,\pi_{{\bf a}}]$.
Notably, any set of $N$ spin observables is $\mathrm{K}$-copy jointly measurable whenever $\mathrm{K}\geq N$, by simply performing different measurements on different copies of the state. However, the scenario becomes nontrivial when $\mathrm{K}<N$. As shown in [23], the three mutually unbiased spin observables $X,Y,Z$ (corresponding to directions $\hat{x},\hat{y},\hat{z}$) are 2-copy jointly measurable for sharpness values up to $\lambda=\sqrt{3}/2$, whereas a naive strategy achieves joint measurability only up to $\lambda=1/\sqrt{2}$ (see Fig.1). We now turn to the anti-parallel ($\uparrow\hskip-1.42271pt\downarrow$) configuration, where the measurement is performed on $\rho_{\vec{m}}\otimes\mathrm{F}(\rho_{\vec{m}})\equiv\rho_{\vec{m}}\otimes\rho_{-\vec{m}}$, with $\rho_{-\vec{m}}:=\tfrac{1}{2}(\mathbf{1}-\vec{m}\cdot\vec{\sigma})$. This leads us to the following definition:
Definition 3.
A set $S$ of $N$ spin observables is jointly measurable on 2-copy $\uparrow\hskip-1.42271pt\downarrow$ configuration if there exists a POVM $\mathcal{G}^{{}_{\uparrow\hskip-1.42271pt\downarrow}}\equiv\{\pi^{{}_{\uparrow\hskip-1.42271pt\downarrow}}_{{\bf a}}\in\mathcal{P}((\mathbb{C}^{2})^{\otimes 2})\penalty 10000\ |\penalty 10000\ \sum_{{\bf a}}\pi^{{}_{\uparrow\hskip-1.42271pt\downarrow}}_{{\bf a}}=\mathbf{1}^{\otimes 2}\}$ such that, for all states $\rho_{\vec{m}}$ and for all $r$, $\mathrm{Tr}[\rho_{\vec{m}}\,\mathrm{P}^{a_{r}}_{\hat{n}_{r}}(\lambda)]=\sum_{{\bf a}\setminus a_{r}}\mathrm{Tr}[(\rho_{\vec{m}}\otimes\rho_{-\vec{m}})\,\pi^{{}_{\uparrow\hskip-1.42271pt\downarrow}}_{{\bf a}}]$.
In what follows, we present a necessary and sufficient condition for joint measurability of spin observables in the $\uparrow\hskip-1.42271pt\downarrow$-configuration.
(a) Naive strategy
(b) Strategy of Carmeli et al.
(c) Antiparallel configuration
Figure 1: (Color online) (a) Naive strategy: Observable $X$ is measured on one copy with red and blue bars respectively denoting the probability of outcomes $1\penalty 10000\ \&\penalty 10000\ -1$ on the given state. The other two observables are jointly measured on the remaining copy with sharpness value $\lambda=1/\sqrt{2}$ [30], with blue (red) part in the red (blue) bar indicating the ‘unsharpness’ in the corresponding outcome. (b) Parallel configuration: All three observables can be jointly measured on two-copies with sharpness value $\lambda=\sqrt{3}/2$ [23] (b) Antiparallel configuration: All three observables can be jointly measured with sharpness value $\lambda=1$ [Theorem 1]
Proposition 1.
A set of spin observables $S=\{\sigma_{\hat{n}_{r}}(\lambda)\}_{r}$ is jointly measurable on 2-copy $\uparrow\hskip-1.42271pt\downarrow$ configuration if and only if there exists a POVM $\mathcal{G}^{{}_{\uparrow\hskip-1.42271pt\downarrow}}\equiv\{\pi^{{}_{\uparrow\hskip-1.42271pt\downarrow}}_{{\bf a}}\penalty 10000\ |\penalty 10000\ \sum_{{\bf a}}\pi^{{}_{\uparrow\hskip-1.42271pt\downarrow}}_{{\bf a}}=\mathbf{1}^{\otimes 2}\}$, such that $\mathbb{P}^{a_{r}}_{\hat{n}_{r}}(\lambda)=\sum_{{\bf a}\setminus a_{r}}\pi^{{}_{\uparrow\hskip-1.42271pt\downarrow}}_{{\bf a}},\penalty 10000\ \forall\penalty 10000\ r,\penalty 10000\ a_{r}$; where $\mathbb{P}^{a_{r}}_{\hat{n}_{r}}(\lambda):=\tfrac{1}{2}(\mathrm{P}^{a_{r}}_{\hat{n}_{r}}(\lambda)\otimes\mathbf{1}+\mathbf{1}\otimes\mathrm{F}[\mathrm{P}^{a_{r}}_{\hat{n}_{r}}(\lambda)])$.
Deferring the proof of Proposition 1 to Supplemental material , here we demonstrate joint measurability of two important sets of spin observables.
Theorem 1.
The set $\mathrm{MUB}\equiv\{X,Y,Z\}$ of three spin observables is jointly measurable on 2-copy $\uparrow\hskip-1.42271pt\downarrow$ configuration.
Proof.
Define the set of operators $\{\Pi^{{}_{\uparrow\hskip-1.42271pt\downarrow}}_{[a,b,c]}\penalty 10000\ |\penalty 10000\ a,b,c=\pm 1\}\subset\mathcal{L}(\mathbb{C}^{2}\otimes\mathbb{C}^{2})$ as
$\Pi^{{}_{\uparrow\hskip-1.42271pt\downarrow}}_{[a,b,c]}$ $:=\tfrac{1}{16}\Bigl(2\mathbf{1}^{\otimes 2}+a[\hskip-1.42271pt[X,\mathbf{1}\penalty 10000\ ]\hskip-1.42271pt]+b[\hskip-1.42271pt[Y,\mathbf{1}\penalty 10000\ ]\hskip-1.42271pt]+c[\hskip-1.42271pt[Z,\mathbf{1}\penalty 10000\ ]\hskip-1.42271pt]$
$-ab\{\hskip-3.41418pt\{X,Y\penalty 10000\ \}\hskip-3.41418pt\}-bc\{\hskip-3.41418pt\{Y,Z\penalty 10000\ \}\hskip-3.41418pt\}-ca\{\hskip-3.41418pt\{Z,X\penalty 10000\ \}\hskip-3.41418pt\}\Bigr),$ (1)
where $\hskip-1.42271pt[U,V\penalty 10000\ ]\hskip-1.42271pt]:=U\otimes V-V\otimes U$ and $\{\hskip-3.41418pt\{U,V\penalty 10000\ \}\hskip-3.41418pt\}:=U\otimes V+V\otimes U$. Direct computation confirms that each $\Pi_{[a,b,c]}^{{}_{\uparrow\hskip-1.42271pt\downarrow}}$ is positive and that $\sum_{a,b,c}\Pi_{[a,b,c]}^{{}_{\uparrow\hskip-1.42271pt\downarrow}}=\mathbf{1}^{\otimes 2}$, hence forming a valid POVM $\mathcal{G}^{\uparrow\hskip-1.42271pt\downarrow}_{\mathrm{MUB}}\equiv\{\Pi_{[a,b,c]}^{{}_{\uparrow\hskip-1.42271pt\downarrow}}\}$. Now, summing over $b$ and $c$ gives $\sum_{b,c}\Pi_{[a,b,c]}^{{}_{\uparrow\hskip-1.42271pt\downarrow}}=\tfrac{1}{4}\left(2\mathbf{1}^{\otimes 2}+a\,[\hskip-1.42271pt[X,\mathbf{1}\right]\hskip-1.42271pt])=\mathbb{P}^{a}_{\hat{x}}$. Similarly, summing over $(c,a)$ and $(a,b)$ yields $\mathbb{P}^{b}_{\hat{y}}$ and $\mathbb{P}^{c}_{\hat{z}}$, respectively. Thus, $\mathcal{G}^{\uparrow\hskip-1.42271pt\downarrow}_{\mathrm{MUB}}$ satisfies the joint measurability condition of Proposition 1, thereby completing the proof. ∎
It is worth noting that the relative angles between any two observables in the set $\{X,Y,Z\}$ are identical. In the Supplemental Material , we present a detailed analysis showing how this symmetry guides construction of the explicit form of the joint POVM of Eq.(1). There, we also identify a class of other symmetric spin triples that are perfectly jointly measurable in the $2$-copy $\uparrow\hskip-1.42271pt\downarrow$ configuration. In the following, we extend this analysis to examine the joint measurability of four other symmetric spin observables in the $2$-copy $\uparrow\hskip-1.42271pt\downarrow$ configuration.
Theorem 2.
The set $\mathrm{SIC}_{4}\equiv\{\sigma_{\hat{n}_{r}}\}_{r=0}^{3}$ of four qubit observables, along the directions $\hat{n}_{r}:=\frac{(-1)^{\bar{\delta}_{r,0}}}{\sqrt{3}}\big((-1)^{\delta_{r,1}},(-1)^{\delta_{r,2}},(-1)^{\delta_{r,3}}\big)$, is jointly measurable on 2-copy $\uparrow\hskip-1.42271pt\downarrow$ configuration.
Proof.
Consider the set of operators $\{\Pi^{{}_{\uparrow\hskip-1.42271pt\downarrow}}_{[\Gamma,a]}\penalty 10000\ |\penalty 10000\ a=\pm 1,\ \Gamma\in\{X,Y,Z\}\}\subset\mathcal{L}(\mathbb{C}^{2}\otimes\mathbb{C}^{2})$, defined as
$\Pi^{{}_{\uparrow\hskip-1.42271pt\downarrow}}_{[\Gamma,a]}:=\frac{1}{12}\Bigl(2\mathbf{1}^{\otimes 2}-2\Gamma^{\otimes 2}+a\sqrt{3}[\hskip-1.42271pt[\Gamma,\mathbf{1}\penalty 10000\ ]\hskip-1.42271pt]+\sum_{\Omega\in\{\Gamma\}^{\footnotesize\complement}}\Omega^{\otimes 2}\Bigr),$ (2)
[…]
$\sum_{\Gamma}\Pi^{{}_{\uparrow\hskip-1.42271pt\downarrow}}_{[\Gamma,+1]}=\tfrac{1}{2}\Bigl(\mathbf{1}^{\otimes 2}+\tfrac{1}{2\sqrt{3}}\sum_{\Gamma}[\hskip-1.42271pt[\Gamma,\mathbf{1}\penalty 10000\ ]\hskip-1.42271pt]\Bigr)=\mathbb{P}^{+1}_{\hat{n}_{0}}.$ (3)
Similarly, summing over the suitable indices reconstructs all $\mathbb{P}^{a_{r}}_{\hat{n}_{r}}$ for $r=0,\cdots 3,\penalty 10000\ \&\penalty 10000\ a_{r}=\pm 1$, satisfying the condition of Proposition 1; and hence completes the proof. ∎
At this point, it is worth recalling a result by Gisin and Popescu [35], which demonstrates that the $\uparrow\hskip-1.42271pt\downarrow$ spin configuration can encode more classical information than the $\uparrow\hskip-1.42271pt\uparrow$ configuration. Remarkably, our Theorems 1 and 2 show that this advantage extends to the joint measurability of multiple spin observables, thereby revealing a novel operational distinction between these configurations.
Recall that a linear map $\Lambda:\mathcal{L}(\mathbb{C}^{2})\to\mathcal{L}(\mathbb{C}^{2})$ is said to be positive (P) if it maps density operators to density operators, i.e., $\Lambda:\mathcal{D}(\mathbb{C}^{2})\to\mathcal{D}(\mathbb{C}^{2})$. However, physicality further requires complete positivity (CP), meaning that the extended map $\mathrm{id}_{d}\otimes\Lambda$ must also preserve positivity on $\mathcal{D}(\mathbb{C}^{d}\otimes\mathbb{C}^{2})$ for all $d\geq 2$, where $\mathrm{id}_{d}$ is the identity map on $\mathbb{C}^{d}$. Thanks to the Choi–Jamiołkowski isomorphism [26, 27], it suffices to verify complete positivity using $d=2$ for qubit maps. In this work, we focus on trace-preserving (TP) maps, which satisfy $\operatorname{Tr}[\Lambda(A)]=\operatorname{Tr}[A]$ for all $A\in\mathcal{L}(\mathbb{C}^{2})$. The spin-flip map $\mathrm{F}$, which appears in the $\uparrow\hskip-1.42271pt\downarrow$ configuration, is an example of a PTP map that is not CPTP (PnCP). Moreover, it is self dual, namely $\mathrm{F}=\mathrm{F}^{\star}$ [36], and $\mathrm{F}\circ\mathrm{F}=\mathrm{id}_{2}$, where $\circ$ denotes sequential composition. This observation motivates us to go beyond the specific $\uparrow\hskip-1.42271pt\uparrow$ and $\uparrow\hskip-1.42271pt\downarrow$ configurations and explore joint measurability in more general configurations of the form $\rho_{\vec{m}}\otimes\Lambda(\rho_{\vec{m}})$, where $\Lambda$ is either a CPTP or a PTP map. At first glance, the question of joint measurability on a configuration $\rho_{\vec{m}}\otimes\Lambda(\rho_{\vec{m}})$ may seem ill-posed when $\Lambda$ is PnCP, as such maps are not physically implementable. However, the question becomes meaningful from an adversarial perspective: a Referee, who knows the state, prepares the bipartite configuration and hands it to an experimentalist, who is unaware of the preparation but aims to implement a joint measurement. Later, we will illustrate how such configurations can arise naturally in various practical scenarios. Before that, we first establish a general result on joint measurability when such broader settings are considered (proof provided in End Matter).
Theorem 3.
For any set of spin observables, the joint measurability relation on the configuration $\rho_{\vec{m}}\otimes\Lambda(\rho_{\vec{m}})$ coincides with that of the $\uparrow\hskip-1.42271pt\uparrow$ configuration whenever $\Lambda$ is a CPTP map.
Theorem 3 shows that to achieve an advantage in joint measurability over the $\uparrow\hskip-1.42271pt\uparrow$ configuration, one must consider configurations of the form $\rho_{\vec{m}}\otimes\Lambda(\rho_{\vec{m}})$, where $\Lambda$ is PnCP. A pertinent example is provided by the family of maps $\mathrm{F}_{\mu}(\rho_{\vec{m}}):=\tfrac{1}{2}(\mathbf{1}-\mu\,\vec{m}\cdot\vec{\sigma})$, which interpolate between the identity map and the spin-flip map, i.e., $\mathrm{F}_{\mu}=\tfrac{1-\mu}{2}\,\text{id}_{2}+\tfrac{1+\mu}{2}\,\mathrm{F},\quad\mu\in[-1,1]$. These maps are PnCP for $\mu\in(1/3,1]$, but CPTP otherwise [37, 38, 39].
Corollary 1.
The configuration $\rho_{\vec{m}}\otimes\mathrm{F}_{\mu}(\rho_{\vec{m}})$ offers an enhanced joint measurability of unsharp MUB observables compared to the $\uparrow\hskip-1.42271pt\uparrow$ configuration whenever $\mu>(\sqrt{3}-1)$.
Proof is discussed in the End Matter. Corollary 1, however, does not demonstrate an advantage over the $\uparrow\hskip-1.42271pt\uparrow$ configuration in the intermediate regime $\mu\in(1/3,\sqrt{3}-1]$, where $\mathrm{F}_{\mu}$ is PnCP but the joint measurability enhancement remains unproven for MUB observables. Whether an advantage exists in this regime for other sets of observables remains open for further investigation. As we discuss in Supplemental , the enhanced performance observed in the $\uparrow\hskip-1.42271pt\downarrow$ case (and in the configuration considered in Corollary 1) can be better appreciated from the mathematical structure of generalized probabilistic theories (GPTs) [40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52]. In the End Matter we introduce the notion of joint measurability for a set of spin observables restricted to a sub-ensemble of qubit states $\mathcal{D}_{SE}\subset\mathcal{D}(\mathbb{C}^{2})$, which provide an experimental means to validate the advantages established in Theorems 1, 2, and Corollary 1. In the remainder of the manuscript, we explore connections to other foundational concepts and discuss the potential applications.
[…]
Subsequent to VAA’s work, Mermin has pointed out that the provocative title of VAA’s paper suggests—without asserting—that Alice determines simultaneous values of three orthogonal spin components, although she does not [56]. Her measurement yields three candidate outcomes, only one of which—the one corresponding to Bob’s actual choice—matches his result; the others do not correspond to any measurement actually performed (thus invoking an element of counter-factuality). In contrast, as established in our Theorem 1, the POVM $\mathcal{G}^{{}_{\uparrow\hskip-1.42271pt\downarrow}}_{\mathrm{MUB}}$ implements a joint measurement of all three spin observables on two particles prepared in the $\uparrow\hskip-1.42271pt\downarrow$ configuration.
Outcome $\backslash$ Retrodiction $X$ $Y$ $Z$
$\mathcal{G}^{\uparrow\hskip-1.42271pt\downarrow}_{\mathrm{MUB}}$ $\mathbb{M}_{+}$ $\xi_{[+1,+1,+1]}$ 1-1 1-1 1-1
$\xi_{[+1,-1,-1]}$ 1-1 +1+1 +1+1
$\xi_{[-1,+1,-1]}$ +1+1 1-1 +1+1
$\xi_{[-1,-1,+1]}$ +1+1 +1+1 1-1
$\mathbb{M}_{-}$ $\xi_{[-1,+1,+1]}$ +1+1 1-1 1-1
$\xi_{[+1,-1,+1]}$ 1-1 +1+1 1-1
$\xi_{[+1,+1,-1]}$ 1-1 1-1 +1+1
$\xi_{[-1,-1,-1]}$ +1+1 +1+1 +1+1
Table 1: (Color online) To win the mean King’s challenge, Alice performs the measurement $\mathrm{M}_{+}=\{\xi_{[a,b,c]}\penalty 10000\ |\penalty 10000\ abc=+1\}$ on her ancilla and the channel system received from Bob. Upon obtaining the outcome $\xi_{[a,b,c]}$, she retrodicts Bob’s outcomes for $X$, $Y$, and $Z$ as a-a, b-b, and c-c, respectively. When Bob restricts his measurements to $X$ and $Z$, a secure key can be established: outcomes corresponding to $\xi_{[+1,-1,-1]}$ and $\xi_{[-1,-1,+1]}$ (marked in red) are used for eavesdropping detection, while the correlation in the remaining outcomes (marked in green) generates the raw key. Although such perfect correlations are absent in $\mathrm{M}_{+}$ when all three observables are included, a key can still be distilled from the outcomes $\xi_{[+1,+1,+1]}$ and $\xi_{[-1,-1,-1]}$ of the POVM $\mathcal{G}^{{}_{\uparrow\hskip-1.42271pt\downarrow}}_{\mathrm{MUB}}$ (highlighted in yellow), with the other outcomes (highlighted in purple) serving to detect potential interference by Eve.
Building on VAA’s mean King strategy, Bub proposed a novel quantum key distribution (QKD) protocol in which Bob restricts his measurements to two choices (e.g., $X$ and $Z$), and Alice performs the $\mathbb{M}_{+}$ measurement on her ancilla and the channel particle returned by Bob [29] (see also [57]). To detect eavesdropping, Alice publicly announces her retrodicted outcomes whenever the projectors $\xi_{[+1,-1,-1]}$ or $\xi_{[-1,-1,+1]}$ click. Bob then checks whether her retrodiction matches his actual result. If no discrepancies arise—indicating no interference by Eve—they use the remaining outcomes to generate a shared raw key (see Table 1). Unlike the BB84 [58] and E91 [59] protocols, this scheme allows key generation without revealing measurement bases, provided Bob’s choices are limited to two observables. However, when Bob chooses among all three Pauli observables ($X$, $Y$, and $Z$), basis disclosure becomes necessary. Interestingly, as shown in Table 1, the POVM $\mathcal{G}^{{}_{\uparrow\hskip-1.42271pt\downarrow}}_{\mathrm{MUB}}$ supports secure key generation even in this case.
Consider the scenario where Bob performs unsharp spin measurements. As shown in [30], the unsharp observables $\{X(\lambda),Z(\lambda)\}$ are jointly measurable for $\lambda\leq 1/\sqrt{2}$, while the triple $\{X(\lambda),Y(\lambda),Z(\lambda)\}$ is compatible for $\lambda\leq 1/\sqrt{3}$. Being compatibility within $\lambda\in[0,1/\sqrt{2}]$, the original protocol yields zero key rate, whereas the modified scheme employing three unsharp observables has the potential to offer an advantage for $\lambda\in(1/\sqrt{3},1/\sqrt{2}]$. While a detailed analysis of this intuition is left for future work, we now turn to another potential application of measurement compatibility in the $\uparrow\hskip-1.42271pt\downarrow$ configuration.
Estimating unknown measurement devices.– Estimation of unknown quantum resources—such as states, channels, or measurement devices—is essential for the reliable implementation of quantum protocols [60, 61, 62, 63, 64, 65]. When the resource is supplied by an untrusted party, a Verifier seeks to verify its identity as efficiently as possible. Here, we consider the task of estimating an unknown qubit projective measurement $\mathrm{M}_{\hat{n}}\equiv\{\mathrm{P}^{a}_{\hat{n}}\penalty 10000\ |\penalty 10000\ a=\pm 1\}$, where the measurement axis $\hat{n}$ is uniformly distributed over the Bloch sphere. The Verifier treats the device as a black box with trusted outcome labels and aims to estimate $\hat{n}$ with minimal use of the device.
[…]
Discussions.– We have explored novel aspects of quantum incompatibility in multi-copy scenarios, demonstrating that the manifestation of measurement incompatibility depends crucially on whether the system copies are prepared in parallel or antiparallel configurations. In particular, we have shown that three mutually orthogonal spin observables – known to be incompatible in the single – copy setting—can become jointly measurable when probed on antiparallel spin pairs, which highlights a distinctive feature of the antiparallel configuration. We further connected our findings to other foundational concepts and explored their potential applications. Specifically, we established a link with the mean King retrodiction task, and discussed how the observed compatibility structure plays a critical role in a quantum key distribution protocol inspired from it. Additionally, we discussed how the joint measurability of three orthogonal spin observables in the antiparallel setup can enhance the efficiency of estimating an unknown measurement device.
Our study welcomes several questions for future investigation. While Theorems 1 and 2 demonstrate the advantage of the antiparallel configuration for specific choices of spin observables, it remains an open question whether similar advantages persist for more general sets of observables. Furthermore, Corollary 1 establishes an advantage of the configuration $\rho_{\vec{m}}\otimes\Lambda(\rho_{\vec{m}})$ over the parallel configuration for a certain class of PnCP maps. It would be interesting to explore whether this advantage extends to all PnCP maps when considering more general observables. Finally, extending the analysis to higher-dimensional quantum systems may reveal additional structures and potential applications.
Acknowledgements.
Acknowledgement: We gratefully acknowledge insightful discussions with Nicolas Gisin on connection between the POVM $\mathcal{G}^{{}_{\uparrow\hskip-1.42271pt\downarrow}}_{\mathrm{MUB}}$ and the elegant joint measurement. KA acknowledges support from the CSIR project $09/0575(19300)/2024$-EMR-I. SRC acknowledges support from University Grants Commission, India (Reference no. 211610113404). SGN acknowledges support from the CSIR project $09/0575(15951)/2022$-EMR-I. MB acknowledge the financial support through the National Quantum Mission (NQM) of the Department of Science and Technology, Government of India.
SUPPLEMENTAL MATERIAL
I Proof of Proposition 1
So far we have analyzed joint measurability for unsharp spin measurements only. However, the notion can be generalized for more general POVMs. A ‘$K$’ outcome qubit POVM $\mathcal{A}$ is the collection of $K\penalty 10000\ (<\infty)$ effects adding up to identity operator. Formally
$\mathcal{A}\equiv\left\{\mathcal{A}^{r}\in\mathcal{E}(\mathbb{C}^{2})\penalty 10000\ |\penalty 10000\ \sum_{r=1}^{K}\mathcal{A}^{r}=\mathbf{1}\right\}.$ (6)
Definition 1S.
A collection of $N$ qubit POVMs $\{\mathcal{A}_{l}\equiv\{\mathcal{A}_{l}^{r_{l}}\}_{r_{l}=1}^{K_{l}}\}_{l=1}^{N}$ is said to be jointly measurable on the configuration $\text{id}_{2}\otimes\Lambda(\rho_{m}\otimes\rho_{m})$, with $\Lambda$ being a PTP map, if there exists a POVM $\mathcal{G}\equiv\left\{\Pi_{\bf r}\in\mathcal{E}((\mathbb{C}^{2})^{\otimes 2})\penalty 10000\ |\penalty 10000\ \sum_{\bf r}\Pi_{\bf r}=\mathbf{1}^{\otimes 2}\right\}\penalty 10000\ \text{with}\penalty 10000\ {\bf r}:=[r_{1},\cdots,r_{N}]\in K_{1}\times\cdots\times K_{N}$, such that for all $\rho_{\vec{m}}\in\mathcal{D}(\mathbb{C}^{2})$
$\sum_{{\bf r}\setminus r_{l}}\operatorname{Tr}\left[\Pi_{\bf r}\left(\rho_{\vec{m}}\otimes\Lambda(\rho_{\vec{m}})\right)\right]=\operatorname{Tr}\left[\mathcal{A}_{l}^{r_{l}}\rho_{\vec{m}}\right],\penalty 10000\ \penalty 10000\ \forall\penalty 10000\ l\in\{1,\cdots,N\}\penalty 10000\ \&\penalty 10000\ r_{l}\in K_{l}.$ (7)
The Proposition 1 in main manuscript consider only unsharp spin observables and $\uparrow\hskip-1.42271pt\downarrow$ configuration. Here, we establish the result for more general POVMs.
Proposition 1S.
A collection of $N$ qubit POVMs $\{\mathcal{A}_{l}\equiv\{\mathcal{A}_{l}^{r_{l}}\}_{r_{l}=1}^{K_{l}}\}_{l=1}^{N}$ is jointly measurable on the state configuration $\rho_{\vec{m}}\otimes\mathrm{F}(\rho_{\vec{m}})$ if and only if there exists a POVM $\mathcal{G}\equiv\{\Pi_{{\bf r}}\in\mathcal{E}((\mathbb{C}^{2})^{\otimes 2})\penalty 10000\ |\penalty 10000\ \sum_{{\bf r}}\Pi_{{\bf r}}=\mathbf{1}^{\otimes 2}\}$, with ${\bf r}:=[r_{1},\cdots,r_{N}]\in K_{1}\times\cdots\times K_{N}$, such that $\mathbb{A}^{r_{l}}_{l}=\sum_{{\bf r}\setminus r_{l}}\Pi_{{\bf r}},\penalty 10000\ \forall\penalty 10000\ l,\penalty 10000\ r_{l}$; where $\mathbb{A}^{r_{l}}_{l}:=\frac{1}{2}\left(\mathcal{A}^{r_{l}}_{l}\otimes\mathbf{1}+\mathbf{1}\otimes\mathrm{F}[\mathcal{A}^{r_{l}}_{l}]\right)$.
Proof.
The proof follows a similar structure as of Theorem 1 in Ref.[23].
Sufficiency
Let there exists a POVM $\mathcal{G}\equiv\{\Pi_{{\bf r}}\in\mathcal{E}((\mathbb{C}^{2})^{\otimes 2})\penalty 10000\ |\penalty 10000\ \sum_{{\bf r}}\Pi_{{\bf r}}=\mathbf{1}^{\otimes 2}\}$ such that
$\sum_{{\bf r}\setminus r_{l}}\Pi_{{\bf r}}=\mathbb{A}^{a_{l}}_{l}=\frac{1}{2}\left(\mathcal{A}^{r_{l}}_{l}\otimes\mathbf{1}+\mathbf{1}\otimes\mathrm{F}[\mathcal{A}^{r_{l}}_{l}]\right),\penalty 10000\ \forall\penalty 10000\ l,\penalty 10000\ r_{l}\penalty 10000\ .$ (8)
Since $\text{id}_{2}$ and $\mathrm{F}$ are self dual, and $\mathrm{F}\circ\mathrm{F}=\text{id}_{2}$, we thus have
$\operatorname{Tr}\Big[\left(\rho_{\vec{m}}\otimes\mathrm{F}(\rho_{\vec{m}})\right)\sum_{{\bf r}\setminus r_{l}}\Pi_{{\bf r}}\Big]$ $=\operatorname{Tr}\left[\left(\rho_{\vec{m}}\otimes\mathrm{F}(\rho_{\vec{m}})\right)\mathbb{A}_{l}^{r_{l}}\right]$
$=\frac{1}{2}\operatorname{Tr}\left[\left(\rho_{\vec{m}}\otimes\mathrm{F}(\rho_{\vec{m}})\right)\text{id}_{2}\otimes\mathrm{F}\left(\mathbf{1}\otimes\mathcal{A}_{l}^{r_{l}}+\mathcal{A}_{l}^{r_{l}}\otimes\mathbf{1}\right)\right]$
$=\frac{1}{2}\operatorname{Tr}\left[\left(\rho_{\vec{m}}\otimes\mathrm{F}\circ\mathrm{F}(\rho_{\vec{m}})\right)\left(\mathbf{1}\otimes\mathcal{A}_{l}^{r_{l}}+\mathcal{A}_{l}^{r_{l}}\otimes\mathbf{1}\right)\right]$
$=\frac{1}{2}\operatorname{Tr}\left[(\rho_{\vec{m}}\otimes\rho_{\vec{m}})\left(\mathbf{1}\otimes\mathcal{A}_{l}^{r_{l}}+\mathcal{A}_{l}^{r_{l}}\otimes\mathbf{1}\right)\right]=\operatorname{Tr}\left[\rho_{\vec{m}}\penalty 10000\ \mathcal{A}_{l}^{r_{l}}\right].$ (9)
This implies that the POVMs $\{\mathcal{A}_{l}\equiv\{\mathcal{A}_{l}^{r_{l}}\}_{r_{l}=1}^{K_{l}}\}_{l=1}^{N}$ are jointly measurable on the state configuration $\rho_{\vec{m}}\otimes\mathrm{F}(\rho_{\vec{m}})$.
Necessity
Let there exists a POVM $\mathcal{G}^{\prime}\equiv\{\Pi^{\prime}_{{\bf r}}\in\mathcal{E}((\mathbb{C}^{2})^{\otimes 2})\penalty 10000\ |\penalty 10000\ \sum_{{\bf r}}\Pi^{\prime}_{{\bf r}}=\mathbf{1}^{\otimes 2}\}$ such that
$\operatorname{Tr}\Big[\left(\rho_{\vec{m}}\otimes\mathrm{F}(\rho_{\vec{m}})\right)\sum_{{\bf r}\setminus r_{l}}\Pi^{\prime}_{{\bf r}}\Big]=\operatorname{Tr}\left[\rho_{\vec{m}}\penalty 10000\ \mathcal{A}_{l}^{r_{l}}\right],\penalty 10000\ \forall\penalty 10000\ l,\penalty 10000\ r_{l},\penalty 10000\ \rho_{\vec{m}}.$ (10)
$\mathrm{F}$ being a completely co-positive map $\mathrm{F}\otimes\mathrm{F}$ becomes positive [70]. Furthermore, since composition of two positive maps is positive, thus $(\mathrm{F}\otimes\mathrm{F})\circ\text{Swap}$ is a PTP map. Therefore it implies existence of another POVM
$\mathcal{G}"\equiv\{(\mathrm{F}\otimes\mathrm{F})\circ\text{Swap}(\Pi^{\prime}_{{\bf r}})$ $:=\Pi"_{{\bf r}}\in\mathcal{E}((\mathbb{C}^{2})^{\otimes 2})\penalty 10000\ |\penalty 10000\ \sum_{{\bf r}}\Pi"_{{\bf r}}=\mathbf{1}^{\otimes 2}\},\penalty 10000\ \penalty 10000\ \text{satisfying}$ (11a)
$\operatorname{Tr}\Big[\left(\rho_{\vec{m}}\otimes\mathrm{F}(\rho_{\vec{m}})\right)\sum_{{\bf r}\setminus r_{l}}\Pi"_{{\bf r}}\Big]$ $=\operatorname{Tr}\Big[\left(\rho_{\vec{m}}\otimes\mathrm{F}(\rho_{\vec{m}})\right)(\mathrm{F}\otimes\mathrm{F})\circ\text{Swap}\Big(\sum_{{\bf r}\setminus r_{l}}\Pi^{\prime}_{{\bf r}}\Big)\Big]$
$=\operatorname{Tr}\Big[\left(\mathrm{F}(\rho_{\vec{m}})\otimes\rho_{\vec{m}}\right)\text{Swap}\Big(\sum_{{\bf r}\setminus r_{l}}\Pi^{\prime}_{{\bf r}}\Big)\Big]$
$=\operatorname{Tr}\Big[\left(\rho_{\vec{m}}\otimes\mathrm{F}(\rho_{\vec{m}})\right)\sum_{{\bf r}\setminus r_{l}}\Pi^{\prime}_{{\bf r}}\Big]=\operatorname{Tr}\left[\rho_{\vec{m}}\penalty 10000\ \mathcal{A}_{l}^{r_{l}}\right],\penalty 10000\ \forall\penalty 10000\ l,\penalty 10000\ r_{l},\penalty 10000\ \rho_{\vec{m}}.$ (11b)
Thus $\mathcal{G}$" also ensures joint measurability of the POVMs $\{\mathcal{A}_{l}\}_{l=1}^{N}$. We therefore can construct another POVM $\mathcal{G}$ of the form
$\mathcal{G}\equiv\Big\{\Pi_{\bf r}:=\frac{1}{2}\Big(\Pi^{\prime}_{\bf r}+\Pi"_{\bf r}\Big)\penalty 10000\ |\penalty 10000\ \Pi_{\bf r}\in\mathcal{E}((\mathbb{C}^{2})^{\otimes 2})\penalty 10000\ \&\penalty 10000\ \sum_{\bf r}\Pi_{\bf r}=\mathbf{1}^{\otimes 2}\Big\},\penalty 10000\ \text{satisfying}$ (12)
$\operatorname{Tr}\Big[\left(\rho_{\vec{m}}\otimes\mathrm{F}(\rho_{\vec{m}})\right)\sum_{{\bf r}\setminus r_{l}}\Pi_{{\bf r}}\Big]=\operatorname{Tr}\left[\rho_{\vec{m}}\penalty 10000\ \mathcal{A}_{l}^{r_{l}}\right],\penalty 10000\ \forall\penalty 10000\ l,\penalty 10000\ r_{l},\penalty 10000\ \rho_{\vec{m}}.$ (13)
Consider now a state $\rho\in\mathcal{D}(\mathbb{C}^{2})$ of the form
$\rho=\frac{1}{2}(\mathbf{1}+t\penalty 10000\ \Delta),$ (14)
with $\Delta=\Delta^{\dagger}\penalty 10000\ \&\penalty 10000\ \operatorname{Tr}[\Delta]=0$ and $|t|\leq 1/\epsilon_{\max}$, where $\epsilon_{\max}$ is the maximum eigenvalue of $\Delta$. Now, using the fact $\mathrm{F}(\Delta)=-\Delta$ and substituting Eq.(14) in Eq.(13) we have
$\operatorname{Tr}\Big[\left(\mathbf{1}^{\otimes 2}+t(\Delta\otimes\mathbf{1}-\mathbf{1}\otimes\Delta)-t^{2}\Delta^{\otimes 2}\right)\mathbb{A}^{r_{l}}_{l}\Big]=2\Big[\operatorname{Tr}[\mathcal{A}_{l}^{r_{l}}]+t\operatorname{Tr}[\Delta\mathcal{A}_{l}^{r_{l}}]\Big],$ (15)
where $\mathbb{A}^{r_{l}}_{l}:=\sum_{{\bf r}\setminus r_{l}}\Pi_{{\bf r}}$. Comparing the coefficients of the same degree in $t$, we get
$\operatorname{Tr}[\mathbb{A}^{r_{l}}_{l}]$ $=2\operatorname{Tr}[\mathcal{A}_{l}^{r_{l}}],$ (16a)
$\operatorname{Tr}[(\Delta\otimes\mathbf{1}-\mathbf{1}\otimes\Delta)\mathbb{A}^{r_{l}}_{l}]$ $=2\operatorname{Tr}[\Delta\mathcal{A}_{l}^{r_{l}}],$ (16b)
$\operatorname{Tr}[\Delta^{\otimes 2}\mathbb{A}^{r_{l}}_{l}]$ $=0.$ (16c)
Since $\Delta$ is Hermitian and traceless, it can be expressed as $\Delta=\sum_{i}a_{i}\sigma_{i}$ with $a_{i}\in\mathbb{R},\penalty 10000\ \text{for}\penalty 10000\ i\in\{x,y,z\}$, and $\sigma$’s being the Pauli operators. Now, substituting the $\Delta=\sum_{i}a_{i}\sigma_{i}$ in Eqs.(16c) & (16b) and comparing same degree coefficients on both sides we respectively obtain
$\operatorname{Tr}[(\sigma_{i}\otimes\sigma_{j})\mathbb{A}^{r_{l}}_{l}]$ $=0,\penalty 10000\ \forall\penalty 10000\ i,j\in\{x,y,z\},$ (17a)
$\operatorname{Tr}[(\sigma_{i}\otimes\mathbf{1}-\mathbf{1}\otimes\sigma_{i})\mathbb{A}^{r_{l}}_{l}]$ $=2\operatorname{Tr}[\sigma_{i}\mathcal{A}_{l}^{r_{l}}],\penalty 10000\ \forall\penalty 10000\ i\in\{x,y,z\}.$ (17b)
Given a particular (but arbitrary) choice of $l$ and $r_{l}$, the operator $\mathbb{A}^{r_{l}}_{l}$ can be expressed in Hilbert-Schmidt form as
$\mathbb{A}^{r_{l}}_{l}=b\mathbf{1}^{\otimes 2}+\sum_{i}c_{i}\sigma_{i}\otimes\mathbf{1}+\sum_{i}d_{i}\mathbf{1}\otimes\sigma_{i}+\sum_{i,j}e_{ij}\sigma_{i}\otimes\sigma_{j}.$ (18)
Since, $\Pi_{\bf r}=(\mathrm{F}\otimes\mathrm{F})\circ\text{Swap}(\Pi_{\bf r})\penalty 10000\ \forall\penalty 10000\ {\bf r}$, we thus have $\mathbb{A}^{r_{l}}_{l}=(\mathrm{F}\otimes\mathrm{F})\circ\text{Swap}\left(\mathbb{A}^{r_{l}}_{l}\right)$, implying $c_{i}=-d_{i}$. Now, from Eqs.(16a), (17a), (17b), & (18) we obtain
$b=\frac{1}{2}\operatorname{Tr}[\mathcal{A}_{l}^{r_{l}}],\penalty 10000\ \penalty 10000\ c_{i}=\frac{1}{4}\operatorname{Tr}[\sigma_{i}\mathcal{A}_{l}^{r_{l}}],\penalty 10000\ \penalty 10000\ e_{ij}=0,\penalty 10000\ \forall\penalty 10000\ i,j\in\{x,y,z\}.$ (19)
Thus we have
$\mathbb{A}^{r_{l}}_{l}$ $=\frac{1}{2}\operatorname{Tr}[\mathcal{A}_{l}^{r_{l}}]\mathbf{1}^{\otimes 2}+\sum_{i}\frac{1}{4}\operatorname{Tr}[\sigma_{i}\mathcal{A}_{l}^{r_{l}}]\sigma_{i}\otimes\mathbf{1}-\sum_{i}\frac{1}{4}\operatorname{Tr}[\sigma_{i}\mathcal{A}_{l}^{r_{l}}]\mathbf{1}\otimes\sigma_{i},$
$=\frac{1}{2}\big(\frac{1}{2}\operatorname{Tr}[\mathcal{A}_{l}^{r_{l}}]\mathbf{1}+\sum_{i}\frac{1}{2}\operatorname{Tr}[\sigma_{i}\mathcal{A}_{l}^{r_{l}}]\sigma_{i}\big)\otimes\mathbf{1}+\frac{1}{2}\mathbf{1}\otimes\big(\frac{1}{2}\operatorname{Tr}[\mathcal{A}_{l}^{r_{l}}]\mathbf{1}-\sum_{i}\frac{1}{2}\operatorname{Tr}[\sigma_{i}\mathcal{A}_{l}^{r_{l}}]\sigma_{i}\big),$
$=\frac{1}{2}\left(\mathcal{A}_{l}^{r_{l}}\otimes\mathbf{1}+\mathbf{1}\otimes\mathrm{F}\big(\mathcal{A}_{l}^{r_{l}}\big)\right).$ (20)
This completes the proof. ∎
II Enhanced measurement compatibility in GPT
Framework of GPT.– Originally developed in the 1960’s [40, 41, 42, 43] and more recently revitalized [44, 45, 46]—offers valuable mathematical insight. In particular, assuming a quantum description for individual systems, a wide range of mathematically consistent composite system descriptions—bounded by the minimal and maximal tensor products—are possible under the constraints of no-signaling and tomographic locality [47, 48, 49, 50, 51, 52]. Among these, the minimal composition offers deeper insight into the extended range of admissible sharpness values for joint measurability observed in the configurations of Corollary 1 of main manuscript.
$\bullet$ Elementary System: An elementary system of a GPT is specified by the triple $(\Omega,\mathbb{E},\mathbb{T})$. The set $\Omega$ denotes the normalized state space, which is generally assumed to be a convex compact set embedded in some real vector space. Mathematically, often it is easy to work with the set of Unnormalized states which form a convex cone. The set $\mathbb{E}$ corresponds to the effect space, where an effect $e\in\mathbb{E}$ yields probability $e(\omega)$ on the state $\omega\in\Omega$. A measurement $M$ is a collection of effects adding to unit effect, i.e. $M\equiv\{e_{k}\penalty 10000\ |\penalty 10000\ \sum_{k}e_{k}=u\}$, where the unite effect is defined as $u(\omega)=1\penalty 10000\ \forall\penalty 10000\ \omega\in\Omega$. Finally, the set $\mathbb{T}$ denotes the reversible transformations allowed on the system. For instance, a qubit in the language of GPT is described as follows:
$\begin{rcases}\Omega\equiv\mathcal{D}(\mathbb{C}^{2}):=\left\{\rho\in\mathcal{L}(\mathbb{C}^{2})\penalty 10000\ |\penalty 10000\ \rho\geq 0\penalty 10000\ \&\operatorname{Tr}[\rho]=1\right\},\\ \mathbb{E}\equiv\mathcal{E}(\mathbb{C}^{2}):=\left\{\pi\in\mathcal{L}(\mathbb{C}^{2})\penalty 10000\ |\penalty 10000\ 0\leq\pi\leq\mathbf{I}_{2}\right\},\\ \mathbb{T}\equiv\mathcal{U}(\mathbb{C}^{2}):=\left\{U\in\mathcal{L}(\mathbb{C}^{2})\penalty 10000\ |\penalty 10000\ UU^{\dagger}=U^{\dagger}U=\mathbf{I}_{2}\right\}\end{rcases},$ (21)
where $\mathcal{U}(\cdot)$ denotes the set of unitary operators.
$\bullet$ Composite System: A GPT also specify the composition rules constrained by no-signaling principle, that prohibits any superluminal communication. Under the assumption of tomographic locality—where global states are fully specified by local measurements [71]—composite systems lie between the minimal and maximal tensor products [47]. For instance, for two elementary qubits, respectively denoted by A and B, the minimal and maximal compositions are defines as
$\begin{rcases}\Omega^{AB}_{\min}&:=\left\{\mathrm{Ch}\{\rho_{{}_{A}}\otimes\rho_{{}_{B}}\}\penalty 10000\ |\penalty 10000\ \rho_{{}_{A}}\in\mathcal{D}(\mathbb{C}^{2}_{A})\penalty 10000\ \&\penalty 10000\ \rho_{{}_{B}}\in\mathcal{D}(\mathbb{C}^{2}_{B})\right\},\\ \mathbb{E}^{AB}_{\min}&:=\left\{\pi_{{}_{AB}}\in\mathcal{L}(\mathbb{C}^{2}_{A}\otimes\mathbb{C}^{2}_{B}),\penalty 10000\ \text{s.t.}\penalty 10000\ 0\leq\operatorname{Tr}[\pi_{{}_{AB}}\rho_{{}_{AB}}]\leq 1,\penalty 10000\ \forall\penalty 10000\ \rho_{{}_{AB}}\in\Omega^{AB}_{\min}\right\},\\ \mathbb{T}^{AB}_{\min}&:=\left\{U_{A}\otimes U_{B}\penalty 10000\ |\penalty 10000\ U_{A}\in\mathcal{U}(\mathbb{C}^{2}_{A}),\penalty 10000\ U_{B}\in\mathcal{U}(\mathbb{C}^{2}_{B})\right\}\end{rcases},$ (22a)
$\begin{rcases}\mathbb{E}^{AB}_{\max}&:=\left\{\mathrm{Ch}\{\pi_{{}_{A}}\otimes\pi_{{}_{B}}\}\penalty 10000\ |\penalty 10000\ \pi_{{}_{A}}\in\mathcal{E}(\mathbb{C}^{2}_{A})\penalty 10000\ \&\penalty 10000\ \pi_{{}_{B}}\in\mathcal{E}(\mathbb{C}^{2}_{B})\right\},\\ \Omega^{AB}_{\max}&:=\left\{\rho_{{}_{AB}}\in\mathcal{L}(\mathbb{C}^{2}_{A}\otimes\mathbb{C}^{2}_{B}),\penalty 10000\ \text{s.t.}\penalty 10000\ 0\leq\operatorname{Tr}[\pi_{{}_{AB}}\rho_{{}_{AB}}\leq 1,\penalty 10000\ \forall\penalty 10000\ \pi_{{}_{AB}}\in\mathbb{E}^{AB}\right\},\penalty 10000\penalty 10000\penalty 10000\\ \mathbb{T}^{AB}_{\max}&:=\left\{U_{A}\otimes U_{B}\penalty 10000\ |\penalty 10000\ U_{A}\in\mathcal{U}(\mathbb{C}^{2}_{A}),\penalty 10000\ U_{B}\in\mathcal{U}(\mathbb{C}^{2}_{B})\right\}\end{rcases},$ (22b)
where $\mathrm{Ch}\{\cdot\}$ denotes convex hull of a set. While minimal composition allows only separable states, the maximal one allows effects that are separable only. The standard quantum composition strictly lies in between,
$\begin{rcases}\Omega^{AB}_{Q}:=\mathcal{D}(\mathbb{C}^{2}_{A}\otimes\mathbb{C}^{2}_{B})\equiv\left\{\rho_{{}_{AB}}\in\mathcal{L}(\mathbb{C}^{2}_{A}\otimes\mathbb{C}^{2}_{B}),\penalty 10000\ \text{s.t.}\penalty 10000\ \rho_{{}_{AB}}\geq 0\penalty 10000\ \&\operatorname{Tr}[\rho_{{}_{AB}}]=1\right\},\\ \mathbb{E}^{AB}_{Q}:=\mathcal{E}(\mathbb{C}^{2}_{A}\otimes\mathbb{C}^{2}_{B})\equiv\left\{\pi_{{}_{AB}}\in\mathcal{L}(\mathbb{C}^{2}_{A}\otimes\mathbb{C}^{2}_{B})\penalty 10000\ |\penalty 10000\ 0\leq\pi_{{}_{AB}}\leq\mathbf{I}_{4}\right\},\\ \mathbb{T}^{AB}_{Q}:=\mathcal{U}(\mathbb{C}^{2}_{A}\otimes\mathbb{C}^{2}_{B})\equiv\left\{U_{{}_{AB}}\in\mathcal{L}(\mathbb{C}^{2}_{A}\otimes\mathbb{C}^{2}_{B})\penalty 10000\ |\penalty 10000\ U_{{}_{AB}}U^{\dagger}_{{}_{AB}}=U^{\dagger}_{{}_{AB}}U_{{}_{AB}}=\mathbf{I}_{4}\right\}\end{rcases}.$ (23)
Notably, maximal Composition includes bipartite states that are not permitted in quantum theory, while minimal Composition allows bipartite effects that are not allowed in quantum composition.
$\bullet$ Measurement compatibility in minimal composition: As noted above, although the state space of minimal composition is restricted, the effect space is enlarged compared to standard quantum theory. This leads to enhanced compatibility for certain measurements, as shown below.
Theorem 1S.
The observables $X^{\lambda},Y^{\lambda},$ and $Z^{\lambda}$ are 2-copy compatible for all $\lambda\in[0,1]$ in the minimal tensor product GPT.
Proof.
Consider the set of operators $\{\Pi^{\#}_{{}^{[a,b,c]}}\penalty 10000\ |\penalty 10000\ a,b,c=\pm 1\}\subset\mathcal{L}(\mathbb{C}^{2}\otimes\mathbb{C}^{2})$, defined as
$\Pi^{\#}_{{}^{[a,b,c]}}$ $:=\frac{1}{16}\Big(2\penalty 10000\ \mathbf{1}^{\otimes 2}+a\{\hskip-3.41418pt\{X,\mathbf{1}\penalty 10000\ \}\hskip-3.41418pt\}+b\{\hskip-3.41418pt\{Y,\mathbf{1}\penalty 10000\ \}\hskip-3.41418pt\}+c\{\hskip-3.41418pt\{Z,\mathbf{1}\penalty 10000\ \}\hskip-3.41418pt\}$
$\hskip 56.9055pt+ab\{\hskip-3.41418pt\{X,Y\penalty 10000\ \}\hskip-3.41418pt\}+bc\{\hskip-3.41418pt\{Y,Z\penalty 10000\ \}\hskip-3.41418pt\}+ca\{\hskip-3.41418pt\{Z,X\penalty 10000\ \}\hskip-3.41418pt\}\Big).$ (24)
Although these operators are not positive and thus cannot be valid effects in quantum theory, they add to the identity operator, i.e.
$\sum_{a,b,c}\Pi^{\#}_{{}^{[a,b,c]}}=\mathbf{1}\otimes\mathbf{1}.$ (25)
Furthermore, on any qubit state $\rho_{\vec{m}}\in\mathcal{D}(\mathbb{C}^{2})$ and for $a,b,c=\pm 1$, they satisfy the following identities
$\sum_{b,c=\pm 1}\operatorname{Tr}\left(\Pi^{\#}_{{}^{[a,b,c]}}\rho^{\otimes 2}_{\vec{m}}\right)$ $=\operatorname{Tr}\left(\mathrm{P}^{a}_{\hat{x}}\rho_{\vec{m}}\right),$ (26a)
$\sum_{a,c=\pm 1}\operatorname{Tr}\left(\Pi^{\#}_{{}^{[a,b,c]}}\rho^{\otimes 2}_{\vec{m}}\right)$ $=\operatorname{Tr}\left(\mathrm{P}^{b}_{\hat{y}}\rho_{\vec{m}}\right),$ (26b)
$\sum_{a,b=\pm 1}\operatorname{Tr}\left(\Pi^{\#}_{{}^{[a,b,c]}}\rho^{\otimes 2}_{\vec{m}}\right)$ $=\operatorname{Tr}\left(\mathrm{P}^{c}_{\hat{z}}\rho_{\vec{m}}\right),$ (26c)
thereby reproducing the measurement statistics of all three Pauli observables. We will now show that these operators are valid effects in minimal composition, i.e, $\Pi^{\#}_{{}^{[a,b,c]}}$ will yield positive probabilities on all two qubit separable states for all $a,b,c$. To establish this it is sufficient to show that it yields positive probabilities on all product states of the form $\rho_{\vec{r}}\otimes\rho_{\vec{s}}$, where $\vec{r}$ and $\vec{s}$ denote the respective Block vectors.
The spin-flip map $\mathrm{F}:\mathcal{L}(\mathbb{C}^{2})\mapsto\mathcal{L}(\mathbb{C}^{2})$ can be written as $\mathrm{F}=T\circ\sigma_{y}$, where $\sigma_{y}$ is the unitary map corresponding to Pauli $Y$ and $T$ is the transpose map in computational basis. From the fact that both $\sigma_{y}$ and $T$ is inner product preserving it follows that
$\operatorname{Tr}(A^{\dagger}B)=\operatorname{Tr}(\mathrm{F}(A^{\dagger})\mathrm{F}(B))),\penalty 10000\ \forall\penalty 10000\ A,B\in\mathcal{L}(\mathcal{H}).$ (27)
One can readily check that
$\Pi^{\#}_{[a,b,c]}=(\mathbf{1}\otimes\mathrm{F})(\Pi^{\uparrow\downarrow}_{[a,b,c]})\penalty 10000\ \forall\penalty 10000\ a,b,c.$ (28)
Since, $\Pi^{\uparrow\downarrow}_{[a,b,c]}$’s are positive we have, for all $a,b,c$, we have
$\operatorname{Tr}\left(\Pi^{\uparrow\downarrow}_{[a,b,c]}\rho_{\vec{r}}\otimes\rho_{\vec{s}}\right)\geq 0,\quad\forall\penalty 10000\ \rho_{\vec{r}},\rho_{\vec{s}}\in\mathcal{D}(\mathbb{C}^{2})$
$\operatorname{Tr}\left(\mathbf{1}\otimes\mathrm{F}(\Pi^{\uparrow\downarrow}_{[a,b,c]})\mathbf{1}\otimes\mathrm{F}(\rho_{\vec{r}}\otimes\rho_{\vec{s}})\right)\geq 0,\quad\forall\penalty 10000\ \rho_{\vec{r}},\rho_{\vec{s}}\in\mathcal{D}(\mathbb{C}^{2})$
$\operatorname{Tr}\left(\Pi^{\#}_{[a,b,c]}\rho_{\vec{r}}\otimes\rho_{-\vec{s}}\right)\geq 0\quad\forall\penalty 10000\ \rho_{\vec{r}},\rho_{\vec{s}}\in\mathcal{D}(\mathbb{C}^{2}),$
This completes the proof. ∎
Naturally the question arises whether a result analogous to Theorem 1S holds for the set of observables $\mathrm{SIC}_{4}\equiv\{\sigma_{\hat{n}_{r}}\}_{r=0}^{3}$ analyzed in Theorem 2 of the main manuscript. We answer this question affirmatively by proving an even more generic result for arbitrary POVM measurements instead of spin observables. Consider $N$ qubit POVMs $\{\mathcal{A}_{l}\equiv\{\mathcal{A}_{l}^{r_{l}}\}_{r_{l}=1}^{K_{l}}\}_{l=1}^{N}$ each consisting of multiple (possibly different) effects. Recalling Definition 1S for antiparallel configuration, the set will be jointly measurable if there exist a POVM
$\mathcal{G}\equiv\left\{\Pi_{\bf r}\in\mathcal{E}((\mathbb{C}^{2})^{\otimes 2})\penalty 10000\ |\penalty 10000\ \sum_{\bf r}\Pi_{\bf r}=\mathbf{1}^{\otimes 2}\right\}\penalty 10000\ \text{with}\penalty 10000\ {\bf r}:=[r_{1},\cdots,r_{N}]\in K_{1}\times\cdots\times K_{N},$ (29)
such that for all $\rho_{\vec{m}}\in\mathcal{D}(\mathbb{C}^{2})$
$\sum_{{\bf r}\setminus r_{l}}\operatorname{Tr}\left[\Pi_{\bf r}\left(\rho_{\vec{m}}\otimes\mathrm{F}(\rho_{\vec{m}})\right)\right]=\operatorname{Tr}\left[\mathcal{A}_{l}^{r_{l}}\rho_{\vec{m}}\right],\penalty 10000\ \penalty 10000\ \forall\penalty 10000\ l\in\{1,\cdots,N\},\penalty 10000\ \&\penalty 10000\ r_{l}\in K_{l}.$ (30)
With this we are now in a position to state our next result.
Theorem 2S.
For any set of qubit POVM measurements $\{\mathcal{A}_{l}\}_{l=1}^{N}$, their joint measurability on the $\uparrow\hskip-1.42271pt\downarrow$ configuration in quantum theory implies their joint measurability on the $\uparrow\hskip-1.42271pt\uparrow$ configuration in minimal tensor product GPT.
Proof.
Let $N$ qubit observables $\{\mathcal{A}_{l}\}_{l=1}^{N}$ be compatible in quantum theory on the $\uparrow\hskip-1.42271pt\downarrow$ configuration. Thus there exists a POVM (29) satisfying Eq.(30). Using the fact that map $\mathrm{F}$ is its own dual we can rewrite Eq.(30) as
$\sum_{{\bf r}\setminus r_{l}}\operatorname{Tr}\left[\text{id}_{2}\otimes\mathrm{F}\left(\Pi_{\bf r}\right)\left(\rho_{\vec{m}}\otimes\rho_{\vec{m}}\right)\right]=\operatorname{Tr}\left[\mathcal{A}_{l}^{r_{l}}\rho_{\vec{m}}\right],\penalty 10000\ \penalty 10000\ \forall\penalty 10000\ l\in\{1,\cdots,N\},\penalty 10000\ r_{l}\in K_{l},\penalty 10000\ \&\penalty 10000\ \rho_{\vec{m}}.$ (31)
In general the operators $\Pi_{\bf r}^{\#}:=\text{id}_{2}\otimes\mathrm{F}\left(\Pi_{\bf r}\right)$ are not positive, but nevertheless they yield positive probabilities on separable states $\sigma_{AB}\in\Omega^{AB}_{\min}$. This follows from the fact that
$\operatorname{Tr}\left[\Pi^{\#}_{\bf r}\sigma_{AB}\right]=\operatorname{Tr}\left[\Pi_{\bf r}\left(\text{id}_{2}\otimes\mathrm{F}(\sigma_{AB})\right)\right]\geq 0,\penalty 10000\ \forall\penalty 10000\ \sigma_{AB}\in\Omega^{AB}_{\min}.$ (32)
Furthermore, $\mathrm{F}$ being unital implies
$\sum_{\bf r}\Pi^{\#}_{\bf r}=\sum_{\bf r}\text{id}_{2}\otimes\mathrm{F}\left(\Pi_{\bf r}\right)=\text{id}_{2}\otimes\mathrm{F}\left(\sum_{\bf r}\Pi_{\bf r}\right)=\text{id}_{2}\otimes\mathrm{F}\left(\mathbf{1}^{\otimes 2}\right)=\mathbf{1}^{\otimes 2}.$ (33)
Thus the collection of operators $\mathcal{A}^{\#}\equiv\left\{\Pi^{\#}_{\bf r}\right\}_{\bf r}$ constitutes a valid measurement in the minimal tensor product GPT. On the other hand, Eq.(31) ensures that this collection of operators also reproduce the statistics of $\{\mathcal{A}_{l}\}_{l=1}^{N}$ on the $\uparrow\hskip-1.42271pt\uparrow$ configuration. This completes the proof. ∎
Remark 1S.
In Theorem 2 of our main manuscript we have shown that the observables $\mathrm{SIC}_{4}$ are compatible in quantum theory on $\uparrow\hskip-1.42271pt\downarrow$ configuration. Now as a corollary of Theorem 2S it follows that these observables are compatible on $\uparrow\hskip-1.42271pt\uparrow$ configuration in the minimal tensor product GPT.
Remark 2S.
One might expect that the converse of Theorem 2S also holds true, namely joint measurability of set of qubit POVMs on $\uparrow\hskip-1.42271pt\uparrow$ configuration in the minimal tensor product GPT would imply their joint measurability on $\uparrow\hskip-1.42271pt\downarrow$ configuration in quantum theory. However, as we argue this is not the case in general. Joint measurability of set of qubit POVMs on $\uparrow\hskip-1.42271pt\uparrow$ configuration in the minimal tensor product GPT implies existence of a collection of operators $\left\{\Pi_{\bf r}\in\mathbb{E}^{AB}_{\min}\penalty 10000\ |\penalty 10000\ \sum_{\bf r}\Pi_{\bf r}=\mathbf{1}^{\otimes 2}\right\}_{\bf r}$, where some of the $\Pi_{\bf r}$’s could be entangled. Consequently, the respective $\text{id}_{2}\otimes\mathrm{F}\left(\Pi_{\bf r}\right)$’s are not positive operators, and hence do not constitute a valid measurement in quantum theory.
III Mean-King problem and Bub’s Cryptography
In the main manuscript we have analyzed the VAA’s strategy for mean King retrodiction task when Bob chooses his measurement from the set $\{X,Y,Z\}$. In this section, we extend the scenario to four observables, where Bob chooses his measurement from the set $\mathrm{SIC}_{4}\equiv\{\sigma_{\hat{n}_{r}}\}_{r=0}^{3}$, where $\hat{n}_{r}:=\frac{(-1)^{\bar{\delta}_{r,0}}}{\sqrt{3}}\big((-1)^{\delta_{r,1}},(-1)^{\delta_{r,2}},(-1)^{\delta_{r,3}}\big)$. Here also Alice shares the channel part of the singlet state with Bob, who sends the updated state back to Alice after performing the measurement of his choice. Alice performs the POVM $\mathcal{G}^{\uparrow\hskip-1.42271pt\downarrow}_{\mathrm{SIC}_{4}}=\{\Pi^{{}_{\uparrow\hskip-1.42271pt\downarrow}}_{[\Gamma,a]}\penalty 10000\ |\penalty 10000\ a=\pm 1,\penalty 10000\ \&\penalty 10000\ \Gamma\in\{X,Y,Z\}\}$, with
$\Pi^{{}_{\uparrow\hskip-1.42271pt\downarrow}}_{[\Gamma,a]}=\frac{1}{12}\Bigl(2\mathbf{1}^{\otimes 2}-2\Gamma^{\otimes 2}+a\sqrt{3}[\hskip-1.42271pt[\Gamma,\mathbf{1}\penalty 10000\ ]\hskip-1.42271pt]+\sum_{\Omega\in\{\Gamma\}^{\footnotesize\complement}}\Omega^{\otimes 2}\Bigr),$ (34)
on the ancilla and channel particle. Depending upon the measurement outcome, she perfectly retrodicts Bob’s measurement outcome as per Table 2.
Outcome $\backslash$ Retrodiction $\sigma_{\hat{n}_{0}}$ $\sigma_{\hat{n}_{1}}$ $\sigma_{\hat{n}_{2}}$ $\sigma_{\hat{n}_{3}}$
$\mathcal{G}^{\uparrow\hskip-1.42271pt\downarrow}_{\mathrm{SIC}_{4}}$ $\Pi^{{}_{\uparrow\hskip-1.42271pt\downarrow}}_{[X,+1]}$ 1-1 1-1 +1+1 +1+1
$\Pi^{{}_{\uparrow\hskip-1.42271pt\downarrow}}_{[X,-1]}$ +1+1 +1+1 1-1 1-1
$\Pi^{{}_{\uparrow\hskip-1.42271pt\downarrow}}_{[Y,+1]}$ 1-1 +1+1 1-1 +1+1
$\Pi^{{}_{\uparrow\hskip-1.42271pt\downarrow}}_{[Y,-1]}$ +1+1 1-1 +1+1 1-1
$\Pi^{{}_{\uparrow\hskip-1.42271pt\downarrow}}_{[Z,+1]}$ 1-1 +1+1 +1+1 1-1
$\Pi^{{}_{\uparrow\hskip-1.42271pt\downarrow}}_{[Z,-1]}$ +1+1 1-1 1-1 +1+1
Table 2: (Color online) Bob performs any one of the measurements from the set $\mathrm{SIC}_{4}\equiv\{\sigma_{\hat{n}_{r}}\}_{r=0}^{3}$, where $\hat{n}_{r}:=\frac{(-1)^{\bar{\delta}_{r,0}}}{\sqrt{3}}\big((-1)^{\delta_{r,1}},(-1)^{\delta_{r,2}},(-1)^{\delta_{r,3}}\big)$. To win the mean King’s challenge, Alice performs the measurement $\mathcal{G}^{\uparrow\hskip-1.42271pt\downarrow}_{\mathrm{SIC}_{4}}=\{\Pi^{{}_{\uparrow\hskip-1.42271pt\downarrow}}_{[\Gamma,a]}\penalty 10000\ |\penalty 10000\ a=\pm 1,\penalty 10000\ \&\penalty 10000\ \Gamma\in\{X,Y,Z\}\}$ on her ancilla and the channel particle (received from Bob). Upon obtaining the outcome, she retrodicts Bob’s outcomes for four observables $\mathrm{SIC}_{4}$ as per the table. When Bob restricts his measurements to $\sigma_{\hat{n}_{0}}$ and $\sigma_{\hat{n}_{1}}$, a secure key can be established: outcomes corresponding to $\Pi^{{}_{\uparrow\hskip-1.42271pt\downarrow}}_{[Y/z,\pm 1]}$ (marked in red) are used for detection of possible eavesdropping, while the correlation in the remaining outcomes (marked in green) generates a raw key.
Bub’s QKD protocol.– The Quantum Key Distribution (QKD) protocol proposed by Bub in also possible in this case if Bob’s measurement choice is limited between any two observables from the set $\mathrm{SIC}_{4}$. For instance, let Bob randomly chooses to perform one of the measurements $\sigma_{\hat{n}_{0}}$, $\sigma_{\hat{n}_{1}}$. Now, if Alice perform the POVM $\mathcal{G}^{\uparrow\hskip-1.42271pt\downarrow}_{\mathrm{SIC}_{4}}$, then they can detect the possible interference of Eve by Alice publicly announcing the outcomes whenever one of the projector $\Pi^{{}_{\uparrow\hskip-1.42271pt\downarrow}}_{[Y/Z,\pm 1]}$ click. In case of no discrepancy they proceed with the QKD protocol, and use the outcomes $\Pi^{{}_{\uparrow\hskip-1.42271pt\downarrow}}_{[X,\pm 1]}$ to generate shared random key (see Table 2).
III.1 Connection between the POVM $\mathcal{G}^{{}_{\uparrow\hskip-1.42271pt\downarrow}}_{\mathrm{MUB}}$ and EJM
All the effects $\Pi^{{}_{\uparrow\hskip-1.42271pt\downarrow}}_{{}^{[a,b,c]}}$’s are rank one, and can be expressed as $\Pi^{{}_{\uparrow\hskip-1.42271pt\downarrow}}_{{}^{[a,b,c]}}\propto\xi_{[a,b,c]}$, where $|\xi:=\ket{\xi}\bra{\xi}$ with
[…]
(a) $\{\xi_{1},\xi_{2},\xi_{4},\xi_{7}\}$
(b) $\{\xi_{l}\}_{l=0}^{7}$
(c) $\{\xi_{0},\xi_{3},\xi_{5},\xi_{6}\}$
Figure 3: (Color online) The vectors $\{\xi_{l}\}_{l=0}^{7}$ (middle). Each vertex is orthogonal to the vertices connected through face diagonal, namely the vectors $\{\xi_{1},\xi_{2},\xi_{4},\xi_{7}\}$ are mutually orthogonal (left) and the vectors $\{\xi_{0},\xi_{3},\xi_{5},\xi_{6}\}$ are mutually orthogonal (right).
[…]
IV Joint measurability of Three Symmetric Spin Observables on Antiparallel configuration
In this section we analyze the joint measurability of a set of three spin observables along directions $\hat{n}_{1},\hat{n}_{2},$ and $\hat{n}_{3}$ that are symmetrically chosen by fixing the relative angles between any two of the spin directions are identical, i.e.,
$\hat{n}_{1}\cdot\hat{n}_{2}=\hat{n}_{2}\cdot\hat{n}_{3}=\hat{n}_{3}\cdot\hat{n}_{1}\penalty 10000\ .$ (37)
Without loss of any generality, by choosing $\hat{t}=\tfrac{1}{\sqrt{3}}(1,1,1)^{\mathrm{T}}$, the above symmetry can be ensured by fixing
$\hat{n}_{1}=(\alpha,\beta,\beta)^{\mathrm{T}},\quad\hat{n}_{2}=(\beta,\alpha,\beta)^{\mathrm{T}},\quad\hat{n}_{3}=(\beta,\beta,\alpha)^{\mathrm{T}},\penalty 10000\ \text{where}$ (38a)
$\alpha:=\tfrac{1}{3}\big(\sqrt{3}\cos\theta+\sqrt{6}\sin\theta\big),\quad\beta:=\tfrac{1}{6}\big(2\sqrt{3}\cos\theta-\sqrt{6}\sin\theta\big),$ (38b)
with $\theta:=\cos^{-1}(\hat{t}\cdot\hat{n}_{i})\in[0,\pi/2]$ (see Fig. 4). The case $\theta=0$ is not interesting as all the observables become identical ($\hat{n}_{1}=\hat{n}_{2}=\hat{n}_{3}=\hat{t}$), and hence are trivially jointly measurable. This symmetric configuration imposes a corresponding symmetry on the parent POVM structure as stated in the following Lemma.
Figure 4: (Color online) The observables $\left\{\sigma_{\hat{n}_{i}}\right\}_{i=1}^{3}$ are symmetrically chosen by fixing the relative angles between any two of them to be identical. Each of them makes an angle $\theta$ with $\hat{t}=\tfrac{1}{\sqrt{3}}(1,1,1)^{\mathrm{T}}$. Inset: view from the top of the $\hat{t}$ vector.
Lemma 1S.
If the symmetric spin observables $\{\sigma_{\hat{n}_{i}}\}_{i=1}^{3}$, with $\hat{n}_{i}$’s defined as in Eqs.(38), are jointly measurable on antiparallel configuration, then there exists a parent POVM $\mathcal{G}^{{}_{\uparrow\hskip-1.42271pt\downarrow,\theta}}\equiv\left\{\Pi^{{}_{\uparrow\hskip-1.42271pt\downarrow,\theta}}_{{[a,b,c]}}:=\Pi^{\theta}_{{[a,b,c]}}\in\mathcal{E}\left(\left(\mathbb{C}^{2}\right)^{\otimes 2}\right)\penalty 10000\ |\penalty 10000\ \sum_{[a,b,c]}\Pi^{\theta}_{[a,b,c]}=\mathbf{1}^{\otimes 2}\right\}$ satisfying the following conditions (for all $a,b,c\in\{\pm 1\}$):
  • (1)
    $\operatorname{Rank}\left[\Pi^{\theta}_{[a,b,c]}\right]=1$ ;
  • (2)
[…]
  • (3)
    $\Pi^{\theta}_{[a,b,c]}=\mathrm{R_{XY}}\otimes\mathrm{R_{XY}}\left(\Pi^{\theta}_{[b,a,c]}\right)=\mathrm{R_{YZ}}\otimes\mathrm{R_{YZ}}\left(\Pi^{\theta}_{[a,c,b]}\right)=\mathrm{R_{ZX}}\otimes\mathrm{R_{ZX}}\left(\Pi^{\theta}_{[c,b,a]}\right)$, where $\mathrm{R_{XY}}$ is the anti unitary operation satisfying $\mathrm{R_{XY}}(\mathbf{1})=\mathbf{1},\penalty 10000\ \mathrm{R_{XY}}(X)=Y,\penalty 10000\ \mathrm{R_{XY}}(Y)=X$ and $\mathrm{R_{XY}}(Z)=Z$. $\mathrm{R_{YZ}}$ and $\mathrm{R_{ZX}}$ are defined analogously.
Proof.
(1) The requirement that $\mathcal{G}^{{}_{\uparrow\hskip-1.42271pt\downarrow,\theta}}$ should reproduces outcome statistics of the observables $\{\sigma_{\hat{n}_{i}}\}_{i=1}^{3}$ demands
$\begin{rcases}&\sum_{b,c}\operatorname{Tr}\left[\Pi^{\theta}_{[a,b,c]}\mathrm{P}^{-a}_{\hat{n}_{1}}\otimes\mathrm{P}^{a}_{\hat{n}_{1}}\right]=0,\penalty 10000\ \forall\penalty 10000\ a;\quad\quad\sum_{a,c}\operatorname{Tr}\left[\Pi^{\theta}_{[a,b,c]}\mathrm{P}^{-b}_{\hat{n}_{2}}\otimes\mathrm{P}^{b}_{\hat{n}_{2}}\right]=0,\penalty 10000\ \forall\penalty 10000\ b;\\ &\hskip 113.81102pt\sum_{a,b}\operatorname{Tr}\left[\Pi^{\theta}_{[a,b,c]}\mathrm{P}^{-c}_{\hat{n}_{3}}\otimes\mathrm{P}^{c}_{\hat{n}_{3}}\right]=0,\penalty 10000\ \forall\penalty 10000\ c;\end{rcases}$ (39)
where $|\mathrm{P}^{k}_{\hat{n}}:=\ket{\psi^{k}_{\hat{n}}}\bra{\psi^{k}_{\hat{n}}}$ be the eigen-projector of the observable $\sigma_{\hat{n}}$ with eigenvalue $k\in\{\pm 1\}$. Eq.(39) further implies that $\forall\penalty 10000\ a,b,c$
$\operatorname{Tr}\left[\Pi^{\theta}_{[a,b,c]}\mathrm{P}^{-a}_{\hat{n}_{1}}\otimes\mathrm{P}^{a}_{\hat{n}_{1}}\right]=0,\penalty 10000\ \penalty 10000\ \operatorname{Tr}\left[\Pi^{\theta}_{[a,b,c]}\mathrm{P}^{-b}_{\hat{n}_{2}}\otimes\mathrm{P}^{b}_{\hat{n}_{2}}\right]=0,\penalty 10000\ \penalty 10000\ \operatorname{Tr}\left[\Pi^{\theta}_{[a,b,c]}\mathrm{P}^{-c}_{\hat{n}_{3}}\otimes\mathrm{P}^{c}_{\hat{n}_{3}}\right]=0.$ (40)
[…]
(2) Let the POVM $\left\{\Pi^{\prime\penalty 10000\ \theta}_{{[a,b,c]}}\right\}$ reproduces measurement statistics of $\left\{\sigma_{\hat{n}_{i}}\right\}_{i=1}^{3}$ for a state $\rho$ on the configuration $\rho\otimes\mathrm{F}(\rho)$. Manifestly, the POVM $\left\{\Pi^{\prime\prime\penalty 10000\ \theta}_{{[a,b,c]}}:=\Pi^{\prime\penalty 10000\ \theta}_{{[-a,-b,-c]}}\right\}$ reproduces measurement statistics of the same observables for a state $\mathrm{F}(\rho)$ on the configuration $\rho\otimes\mathrm{F}(\rho)$, or equivalently $\rho$ on the configuration $\mathrm{F}(\rho)\otimes\rho$. Let
$\Pi^{\theta}_{{[a,b,c]}}:=\frac{1}{2}\Pi^{\prime\penalty 10000\ \theta}_{{[a,b,c]}}+\frac{1}{2}(\mathrm{F}\otimes\mathrm{F})\left(\Pi^{\prime\prime\penalty 10000\ \theta}_{{[a,b,c]}}\right).$ (41)
Applying $\mathrm{F}\otimes\mathrm{F}$ on both sides we get
$(\mathrm{F}\otimes\mathrm{F})\left(\Pi^{\theta}_{{[a,b,c]}}\right)$ $=\frac{1}{2}(\mathrm{F}\otimes\mathrm{F})\left(\Pi^{\prime\penalty 10000\ \theta}_{{[a,b,c]}}\right)+\frac{1}{2}\Pi^{\prime\prime\penalty 10000\ \theta}_{{[a,b,c]}}$
$=\frac{1}{2}(\mathrm{F}\otimes\mathrm{F})\left(\Pi^{\prime\prime\penalty 10000\ \theta}_{{[-a,-b,-c]}}\right)+\frac{1}{2}\Pi^{\prime\penalty 10000\ \theta}_{{[-a,-b,-c]}}=\Pi^{\theta}_{{[-a,-b,-c]}}.$ (42)
(3) Let the POVM $\left\{\Pi^{\prime\penalty 10000\ \theta}_{{[a,b,c]}}\right\}$ reproduces measurement statistics of $\left\{\sigma_{\hat{n}_{i}}\right\}_{i=1}^{3}$ for a state $\rho$ on the configuration $\rho\otimes\mathrm{F}(\rho)$. Manifestly, the POVM $\left\{\Pi^{\prime\prime\penalty 10000\ \theta}_{{[a,b,c]}}:=\Pi^{\prime\penalty 10000\ \theta}_{{[b,a,c]}}\right\}$ reproduces measurement statistics of the same observables for a state $\mathrm{R_{XY}}(\rho)$ on the configuration $\rho\otimes\mathrm{F}(\rho)$, or equivalently $\rho$ on the configuration $(\mathrm{R_{XY}}\otimes\mathrm{R_{XY}})(\rho\otimes\mathrm{F}(\rho))$. Let
$\Pi^{\theta}_{{[a,b,c]}}:=\frac{1}{2}\Pi^{\prime\penalty 10000\ \theta}_{{[a,b,c]}}+\frac{1}{2}(\mathrm{R_{XY}}\otimes\mathrm{R_{XY}})\left(\Pi^{\prime\prime\penalty 10000\ \theta}_{{[a,b,c]}}\right).$ (43)
Applying $\mathrm{R_{XY}}\otimes\mathrm{R_{XY}}$ on both sides we get
$(\mathrm{R_{XY}}\otimes\mathrm{R_{XY}})\left(\Pi^{\theta}_{{[a,b,c]}}\right)$ $=\frac{1}{2}(\mathrm{R_{XY}}\otimes\mathrm{R_{XY}})\left(\Pi^{\prime\penalty 10000\ \theta}_{{[a,b,c]}}\right)+\frac{1}{2}\Pi^{\prime\prime\penalty 10000\ \theta}_{{[a,b,c]}}$
$=\frac{1}{2}(\mathrm{R_{XY}}\otimes\mathrm{R_{XY}})\left(\Pi^{\prime\prime\penalty 10000\ \theta}_{{[b,a,c]}}\right)+\frac{1}{2}\Pi^{\prime\penalty 10000\ \theta}_{{[b,a,c]}}=\Pi^{\theta}_{{[b,a,c]}}.$ (44)
Similarly, $\Pi^{\theta}_{[a,c,b]}=\mathrm{R_{YZ}}\otimes\mathrm{R_{YZ}}\left(\Pi^{\theta}_{[a,b,c]}\right)$and $\Pi^{\theta}_{[c,b,a]}=\mathrm{R_{ZX}}\otimes\mathrm{R_{ZX}}\left(\Pi^{\theta}_{[a,b,c]}\right)$.
This completes the proof of Lemma 1S. ∎
[…]
Lemma 2S.
[…]
$u=\frac{4}{3(\cos 2\theta-1)},\penalty 10000\ \penalty 10000\ \penalty 10000\ v=\frac{-4}{(5+3\cos 2\theta)},\penalty 10000\ \penalty 10000\ \penalty 10000\ w=\frac{-4(7+9\cos 2\theta)}{(5+3\cos 2\theta)^{2}}\penalty 10000\ .$
[…]
Proof.
[…]
$K_{\chi}=\mathbf{1}+\sum_{i=1}^{3}u_{i}\mathrm{P}^{-1}_{\hat{n}_{i}}.$ (45)
[…]
This uniquely solves $u_{i}$’s and yields
$u_{1}=u_{2}=u_{3}:=u=\frac{4}{3(\cos 2\theta-1)}.$ (46)
Similar analysis for $K_{\eta}$ fixes the values of $v$ and $w$, and completes the proof. ∎
We have almost identified the desired POVM apart from the normalization factors appearing with its Rank-1 effects. These normalization factors are dependent on the angle $\theta$ and put constraints on existence of the parent POVM as established in our next Theorem.
Theorem 3S.
The set of observables $\left\{\sigma_{\hat{n}_{i}}\right\}_{i=1}^{3}$ is jointly measurable on 2-copy $\uparrow\hskip-1.42271pt\downarrow$ configuration if and only if $\theta\in[0,\cos^{-1}(1/3)]$.
Proof.
Assuming normalization for $\Pi_{[1,1,1]}$ and $\Pi_{[1,1,-1]}$ as $N$ and $M$ respectively and using Lemma 2S, we get
$\Pi_{[1,1,1]}$ $|\displaystyle=N\ket{\chi}\bra{\chi}$
$=N\Big(p\penalty 10000\ \mathbf{1}^{\otimes 2}+q\penalty 10000\ ([\hskip-1.42271pt[X,\mathbf{1}\penalty 10000\ ]\hskip-1.42271pt]+[\hskip-1.42271pt[Y,\mathbf{1}\penalty 10000\ ]\hskip-1.42271pt]+[\hskip-1.42271pt[Z,\mathbf{1}\penalty 10000\ ]\hskip-1.42271pt])$
$\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ +r\penalty 10000\ (\{\hskip-3.41418pt\{X,Y\penalty 10000\ \}\hskip-3.41418pt\}+\{\hskip-3.41418pt\{Y,Z\penalty 10000\ \}\hskip-3.41418pt\}+\{\hskip-3.41418pt\{Z,X\penalty 10000\ \}\hskip-3.41418pt\})+s\penalty 10000\ \big(X^{\otimes 2}+Y^{\otimes 2}+Z^{\otimes 2}\big)\penalty 10000\ \Big)\penalty 10000\ ,$ (47a)
$\Pi_{[1,1,-1]}$ $|\displaystyle=M\ket{\eta}\bra{\eta}$
$=M\Big(p_{1}\mathbf{1}^{\otimes 2}+q_{1}([\hskip-1.42271pt[X,\mathbf{1}\penalty 10000\ ]\hskip-1.42271pt]+[\hskip-1.42271pt[Y,\mathbf{1}\penalty 10000\ ]\hskip-1.42271pt])\penalty 10000\ +q_{2}\penalty 10000\ [\hskip-1.42271pt[Z,\mathbf{1}\penalty 10000\ ]\hskip-1.42271pt]$
$\penalty 10000\ \penalty 10000\ \penalty 10000\ +r_{1}\penalty 10000\ (\{\hskip-3.41418pt\{Y,Z\penalty 10000\ \}\hskip-3.41418pt\}+\{\hskip-3.41418pt\{Z,X\penalty 10000\ \}\hskip-3.41418pt\})+r_{2}\{\hskip-3.41418pt\{X,Y\penalty 10000\ \}\hskip-3.41418pt\}+s_{1}\big(X^{\otimes 2}+Y^{\otimes 2}\big)\penalty 10000\ +s_{2}Z^{\otimes 2}\penalty 10000\ \Big),$ (47b)
where,
$\begin{rcases}p=1+\frac{3}{2}u\left(2+u\left(2+2\alpha\beta+\beta^{2}\right)\right),\\ q=\frac{1}{2}u(2+3u)(\alpha+2\beta),\\ r=-\frac{1}{2}u^{2}(1+2\beta(2\alpha+\beta)),\\ s=-1+\frac{1}{2}u\left(-6+u\left(-4+2\alpha\beta+\beta^{2}\right)\right)\end{rcases},$ (48a)
$\begin{rcases}p_{1}=\frac{1}{2}\left(2+4v+w(2+w)-2vw\left(-1+2\alpha\beta+\beta^{2}\right)+v^{2}\left(3+2\alpha\beta+\beta^{2}\right)\right),\\ q_{1}=\frac{1}{2}(2+2v+w)(-w\beta+v(\alpha+\beta)),\\ q_{2}=\frac{1}{2}(2+2v+w)(-w\alpha+2v\beta),\\ r_{1}=\frac{1}{2}\left(-w^{2}\alpha\beta-2v^{2}\beta(\alpha+\beta)+v(w+w\alpha\beta)\right),\\ r_{2}=\frac{1}{2}\left(-w^{2}\beta^{2}+2vw\beta(\alpha+\beta)+v^{2}\left(-1-2\alpha\beta+\beta^{2}\right)\right),\\ s_{1}=-(1+v)(1+v+w)-vw\alpha\beta-\frac{1}{2}\left(-2v^{2}+w^{2}\right)\beta^{2},\\ s_{2}=\frac{1}{2}\left(-2+v^{2}\left(-1+2\alpha\beta-3\beta^{2}\right)-2v\left(2+w+w\beta^{2}\right)+w\left(-2+w\left(-1+2\beta^{2}\right)\right)\right)\end{rcases}.$ (48b)
Now, the reproducibility condition of Proposition 1S for the parent POVM further demands
$\begin{rcases}&Np+3Mp_{1}=\frac{1}{2},\penalty 10000\ \penalty 10000\ Nq+M(2q_{1}-q_{2})=\frac{\alpha}{4},\penalty 10000\ \penalty 10000\ Nq+Mq_{2}=\frac{\beta}{4},\\ &\hskip 28.45274ptNr+M(2r_{1}+r_{2})=0,\penalty 10000\ \penalty 10000\ Ns+M(2s_{1}+s_{2})=0.\end{rcases}$ (49)
It turns out that the equations in (49) are linearly dependent and yield unique solution for $N$ and $M$ as below
$N(\theta)$ $=\frac{(7+9\cos 2\theta)\tan^{4}\theta}{128},$ (50a)
$M(\theta)$ $=\frac{(5+3\cos 2\theta)^{4}\tan\theta}{128\penalty 10000\ (3\sin 2\theta)^{3}}.$ (50b)
Note that, $N(\theta)\geq 0$ for $\theta\in(0,\cos^{-1}(1/3)]$ and yield consistent parent POVM. However, $N(\theta)<0$ for $\theta\in(\cos^{-1}(1/3),\pi/2]$ (see Fig. 5), and hence no parent POVM exists. This complete the proof.
Figure 5: (Color online) Plot of the normalization factors $N(\theta)$ and $M(\theta)$ with $\theta$. $N(\theta)<0$ for $\theta\in(\cos^{-1}(1/3),\pi/2]$.
Remark 3S.
Putting $\theta=\cos^{-1}(1/\sqrt{3})$ in Theorem 3S we obtain,
$\begin{rcases}N=M=1/8;\penalty 10000\ \penalty 10000\ p=p_{1}=1;\penalty 10000\ \penalty 10000\ q=q_{1}=r_{1}=1/2;\penalty 10000\ \penalty 10000\ \\ \hskip 28.45274ptr=q_{2}=r_{2}=-1/2;\penalty 10000\ \penalty 10000\ s=s_{1}=s_{2}=0\penalty 10000\ .\end{rcases}$
This is the particular case of observables $\{X,Y,Z\}$ discussed in Theorem 1 of the main manuscript.
Remark 4S.
Putting $\theta=\cos^{-1}(1/3)$ in Theorem 3S we obtain,
$\begin{rcases}N=0;\penalty 10000\ \penalty 10000\ M=1/6;\penalty 10000\ \penalty 10000\ p_{1}=1;\penalty 10000\ \penalty 10000\ q_{1}=1/\sqrt{3};\penalty 10000\ \penalty 10000\ q_{2}=-1/2\sqrt{3};\\ \hskip 28.45274ptr_{1}=s_{2}=1/3;\penalty 10000\ \penalty 10000\ r_{2}=-2/3,\penalty 10000\ \penalty 10000\ s_{1}=-1/6\penalty 10000\ .\end{rcases}$
Accordingly we have $\Pi_{[1,1,1]}=0=\Pi_{[-1,-1,-1]}$, and the other six effects are given by
$\frac{1}{6}\penalty 10000\ \times\penalty 10000\$
$\mathbf{1}^{\otimes 2}$ $\hskip-1.42271pt[X,\mathbf{1}$]] $\hskip-1.42271pt[Y,\mathbf{1}$]] $\hskip-1.42271pt[Z,\mathbf{1}$]] $\{\hskip-3.41418pt\{X,Y$}} $\{\hskip-3.41418pt\{Y,Z$}} $\{\hskip-3.41418pt\{Z,X$}} $X^{\otimes 2}$ $Y^{\otimes 2}$ $Z^{\otimes 2}$
$\Pi_{[1,1,-1]}$ 11 $\tfrac{1}{\sqrt{3}}$ $\tfrac{1}{\sqrt{3}}$ $\tfrac{1}{2\sqrt{3}}$ $\tfrac{2}{3}$ $\tfrac{1}{3}$ $\tfrac{1}{3}$ $\tfrac{1}{6}$ $\tfrac{1}{6}$ $\tfrac{1}{3}$
$\Pi_{[1,-1,1]}$ 11 $\tfrac{1}{\sqrt{3}}$ $\tfrac{1}{2\sqrt{3}}$ $\tfrac{1}{\sqrt{3}}$ $\tfrac{1}{3}$ $\tfrac{1}{3}$ $\tfrac{2}{3}$ $\tfrac{1}{6}$ $\tfrac{1}{3}$ $\tfrac{1}{6}$
$\Pi_{[-1,1,1]}$ 11 $\tfrac{1}{2\sqrt{3}}$ $\tfrac{1}{\sqrt{3}}$ $\tfrac{1}{\sqrt{3}}$ $\tfrac{1}{3}$ $\tfrac{2}{3}$ $\tfrac{1}{3}$ $\tfrac{1}{3}$ $\tfrac{1}{6}$ $\tfrac{1}{6}$
$\Pi_{[-1,-1,1]}$ 11 $\tfrac{1}{\sqrt{3}}$ $\tfrac{1}{\sqrt{3}}$ $\tfrac{1}{2\sqrt{3}}$ $\tfrac{2}{3}$ $\tfrac{1}{3}$ $\tfrac{1}{3}$ $\tfrac{1}{6}$ $\tfrac{1}{6}$ $\tfrac{1}{3}$
$\Pi_{[-1,1,-1]}$ 11 $\tfrac{1}{\sqrt{3}}$ $\tfrac{1}{2\sqrt{3}}$ $\tfrac{1}{\sqrt{3}}$ $\tfrac{1}{3}$ $\tfrac{1}{3}$ $\tfrac{2}{3}$ $\tfrac{1}{6}$ $\tfrac{1}{3}$ $\tfrac{1}{6}$
$\Pi_{[1,-1,-1]}$ 11 $\tfrac{1}{2\sqrt{3}}$ $\tfrac{1}{\sqrt{3}}$ $\tfrac{1}{\sqrt{3}}$ $\tfrac{1}{3}$ $\tfrac{2}{3}$ $\tfrac{1}{3}$ $\tfrac{1}{3}$ $\tfrac{1}{6}$ $\tfrac{1}{6}$
Table 3: The coefficients of the effects are show in the Table. Each values in the table are multiplied by $1/6$, as indicated at the right side of the Table.
Note that,
$\operatorname{Tr}\left[\left(\Pi_{[1,1,-1]}+\Pi_{[1,-1,1]}+\Pi_{[1,-1,-1]}\right)\Big(\rho\otimes\mathrm{F}(\rho)\Big)\right]=\operatorname{Tr}\left[\mathrm{P}^{+1}_{\hat{m}_{1}}\rho\right],$ (51a)
$\operatorname{Tr}\left[\left(\Pi_{[1,1,-1]}+\Pi_{[-1,1,1]}+\Pi_{[-1,1,-1]}\right)\Big(\rho\otimes\mathrm{F}(\rho)\Big)\right]=\operatorname{Tr}\left[\mathrm{P}^{+1}_{\hat{m}_{2}}\rho\right],$ (51b)
$\operatorname{Tr}\left[\left(\Pi_{[1,-1,1]}+\Pi_{[-1,1,1]}+\Pi_{[-1,-1,1]}\right)\Big(\rho\otimes\mathrm{F}(\rho)\Big)\right]=\operatorname{Tr}\left[\mathrm{P}^{+1}_{\hat{m}_{3}}\rho\right],$ (51c)
where
$\hat{m}_{1}=\tfrac{1}{3\sqrt{3}}(5,-1,-1)^{\mathrm{T}},\penalty 10000\ \penalty 10000\ \hat{m}_{2}=\tfrac{1}{3\sqrt{3}}(-1,5,-1)^{\mathrm{T}},\penalty 10000\ \penalty 10000\ \hat{m}_{3}=\tfrac{1}{3\sqrt{3}}(-1,-1,5)^{\mathrm{T}}.$ (52)
Furthermore, we also have
$\operatorname{Tr}\left[\left(\Pi_{[-1,-1,1]}+\Pi_{[-1,1,-1]}+\Pi_{[1,-1,-1]}\right)\Big(\rho\otimes\mathrm{F}(\rho)\Big)\right]=\operatorname{Tr}\left[\mathrm{P}^{+1}_{\hat{m}_{0}}\rho\right],$ (53)
$\hskip 113.81102pt\text{where}\penalty 10000\ \hat{m}_{0}=\tfrac{1}{\sqrt{3}}(-1,-1,-1)^{\mathrm{T}}.$ (54)
Also note that
$\hat{m}_{i}\cdot\hat{m}_{j\neq i}=-\tfrac{1}{3},\penalty 10000\ \forall\penalty 10000\ i,j\in\{0,1,2,3\}\penalty 10000\ .$ (55)
Thus, the set of vectors $\{\hat{m}_{r}\}_{r=0}^{3}$ are related with the vectors $\{\hat{n}_{r}\}_{r=0}^{3}$ in Theorem 2 of the main manuscript as follows
$\hat{m}_{0}=-\hat{n}_{0},\penalty 10000\ \penalty 10000\ \&\penalty 10000\ \penalty 10000\ \hat{m}_{r}=\mathcal{R}^{\hat{m}_{0}}_{\varphi}(-\hat{n}_{r}),\penalty 10000\ \penalty 10000\ \text{for}\penalty 10000\ \penalty 10000\ r\in\{1,2,3\},$ (56)
where $\mathcal{R}^{\hat{m}_{0}}_{\varphi}$ denotes rotation about $\hat{m}_{0}$ by an angle $\varphi=\cos^{-1}(7/9)$.